SubjectCover

Question 13 of 18 in Handling Data 7

Handling Data 7

3 marks · Normal

A box contains 5 pens: 3 blue and 2 black. Two pens are taken at random, one after the other, without replacement.

Work out the probability that both pens are black.

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

0.1 (±0.01)

Also accept: 1/10

Common mistakes: Using with replacement method 0.08

Show mark scheme and worked solution

Marking points

M1 · 1 mark · States probability first pen is black: 2/5.

M2 · 1 mark · States probability second pen is black if first was black: 1/4.

A1 · 1 mark · Multiplies: 2/5 × 1/4 = 2/20 = 1/10 = 0.1.

Worked solution

First pen black: 2/5

Second pen black (one less, total now 4): 1/4

Report a problem with this question
What's wrong?